Quantity demanded= 20000-3P



quantity supplied= 15000+2P



a. calculate the equilibrium price and quantity.



b. calculate the price elasticity of supply using the point method when the economy is in equilibrium.

1
Expert's answer
2021-12-07T10:34:41-0500

a.

At equilibrium Qd = Qs

20,000-3P = 15,000 + 2P

20,000 – 15,000 = 2P+3P

5,000 = 5P

P = 5,000/5 = 1,000

The equilibrium price = $1,000

We will substitute the equilibrium price in either the demand equation or supply equation to get the equilibrium quantity.

Qd = 20,000-3P

But,

P = 1,000

Hence,

Qd = 20,000-3(1,000)

Q = 17,000 units

The equilibrium quantity = 17,000 units

b.

The price elasticity of supply

Price elasticity of supply (PES) = (change in Quantity /change in Price) * (Price / Quantity)

But,

Change in Quantity /change in Price = 2 (obtained by differentiating the supply equation)

Price (P) = 1,000

Quantity (Q) = 17,000

Hence, PES = (2) * (1,000 / 17,000) = 0.12

The price elasticity of supply = 0.12